OpenStudy (anonymous):

approximate the area of the region R which is bounded above b the parbola y=x^2+1, below by the x-axis, on the left by the line x=-2 and on the right by the line x=3. can u find the answer?

OpenStudy (mr.math):

What about an exact answer? I hate approximation!

OpenStudy (turingtest):

can you draw a picture of the graphs?

OpenStudy (anonymous):

subdivide [-2,3] into n=5 subintervals

OpenStudy (turingtest):

well that changes things a bit...

OpenStudy (anonymous):

triangle ix=b-a/n

OpenStudy (earthcitizen):

can you give the equation ?

OpenStudy (turingtest):

\[A\approx \sum_{i=1}^{n}f(x_i)\Delta x\]where\[\Delta x=\frac{b-a}{n}\]

OpenStudy (anonymous):

A_R≅f(∝i)∆ix

OpenStudy (turingtest):

something like that

OpenStudy (earthcitizen):

what rule or formula was that ?

OpenStudy (turingtest):

Reimann sum

OpenStudy (turingtest):

http://en.wikipedia.org/wiki/Riemann_sum

OpenStudy (earthcitizen):

Trapezoidal rule

OpenStudy (turingtest):

They have it written slightly differently because they represent each part of delta x separately. Since Delta represent "change in" though, they are the same. ^Yes that is the same.

OpenStudy (earthcitizen):

\[\int\limits_{-2}^{3} x ^{2}+1\]

OpenStudy (turingtest):

No that will not be an approximation.

OpenStudy (earthcitizen):

tru

OpenStudy (earthcitizen):

nt a fan of approx..lol

OpenStudy (earthcitizen):

how will you approxim8 ?

OpenStudy (turingtest):

It hasn't told us what kind of Reimann sum to do. There are a few, but above is a formula I gave above is for approximation from the right endpoints of the rectangles I think.

OpenStudy (turingtest):

More or less we are breaking it up like this: